OpenStudy (anonymous):

(a-3b)^(2) multiply i need work shown step by step

OpenStudy (bahrom7893):

(x+y)^2=x^2+2xy+y^2

OpenStudy (bahrom7893):

and (x-y)^2=x^2-2xy+y^2

OpenStudy (bahrom7893):

in you case you would have: (a-3b)^2=a^2-2*3ab+9b^2

OpenStudy (bahrom7893):

or a^2-6ab+9b^2